# Intermittent localization and fast spatial learning by non-Markov random walks with decaying memory

### v1

@@ V.2 Time-dependent part
The analysis of the spatial part have allowed us to find the largest eigenvalue (43) of the perturbation spectrum,
@@
In the temporal part, $\lambda$ only appears in the first term of the rhs of Eq. (23) for the evolution of $h(t)$, which involves the kernel $\pi_{t,t^{\prime}}$. We expect the amplitude to grow with time if $\lambda$ is large enough, say, above a threshold value denoted as $\lambda_{c}^{(I)}$, and to decay to 0 if $\lambda<\lambda_{c}^{(I)}$. The largest eigenvalue $\lambda_{max}$ corresponds to the most “dangerous” mode, that dominates the dynamics at late times by growing the fastest or decaying the slowest. Heuristically, these two types of asymptotic behaviors are separated by the marginal case at $\lambda=\lambda_{c}$ where $h(t)\to h^{*}$ at late times, with $h^{*}$ a constant. Substituting $h(t)$ by a constant in Eq. (23) and taking $t\to\infty$, one obtains
@@
For memory kernels of the form (2), where $F(\tau)$ is assumed to decay or to stay constant, the initial times in the interval $t^{\prime}\in[0,t_{0}]$ contribute very little at large $t$ to the distribution $\pi_{t,t^{\prime}}$. Hence the sum in Eq. (46) is equal to unity due to normalization and the equation simplifies to
@@
These conclusions agree with some of our results: in Section IV.1, fully localized states appear to be stable attractors of the dynamics in the simulations, for all memory kernels that are sufficiently long-ranged (see Fig. 1b). This is actually the regime where the de-correlation approximation used in the stability analysis is expected to be most accurate and valid.
Unfortunately, the same conclusion predicts that localization in short-range memory systems should be stable, too, which is inconsistent with the observation of intermittent localization. The theory is thus unable to describe the intermittent localization phenomenon, for which one could have, for instance, at least one unstable mode or $\lambda_{max}>\lambda_{c}$. This suggests that during the evolution of localization toward diffusive states, temporal correlations start to play a crucial role and must be taken into account.
@@
Let us return to the case of the power-law memory kernel (15) with $0\leq\beta\leq 1$, a regime where the above linear stability analysis holds in principle.
@@ VI.1 Relaxation of the slowest mode
How the amplitude of a perturbation relaxes to zero does depend on the memory kernel. We can apply the results of the stability analysis to the study of the relaxation of the leading eigenmode with $\lambda=\lambda_{max}=1-q$, which corresponds to Eq. (40) with $k=0$. At large $t$, we approximate the sums by integrals and time is taken as continuous; in addition h(t+1)h(t)h(t+1)-h(t) is replaced by the time-derivative $\dot{h}(t)$. Equation (23) with $\lambda=1-q$ and $t_{0}=0$ becomes
@@
$h(t)=t^{\nu},$ (50)
for the solution of Eq. (48) at late times, with $\nu$ a negative exponent to be determined. Following a similar method as in [34], we substitute $h(t^{\prime})$ by the form (50) over the whole time domain, choose a positive $\varepsilon\ll 1$ and a large $t$ such that $\varepsilon t\gg 1$. We then make the change $\tau=t-t^{\prime}$ and decompose,
@@
The first and second integrals in the rhs of Eq. (51) are ${\cal O}\left((\varepsilon t)^{1-\beta}\right)$ and ${\cal O}\left(\varepsilon(\varepsilon t)^{1-\beta}\right)$, respectively, and can be neglected compared to the third term, which is ${\cal O}\left(t^{1-\beta}\right)$. To see this, let us replace $(1+\tau)^{\beta}$ by $\tau^{\beta}$ in the denominator of the third term (as $\epsilon t\gg 1$),
@@
We analyze below the different cases.
$\bullet$ Case $\beta=0$ : When memory is perfect, the solution of Eq. (53) is simply $\nu_{\beta=0}=-\gamma P_{0}^{(st)}$, which belongs to the interval (1,0)(-1,0), as expected. Hence,
@@
Hence the amplitude of most dangerous eigenmode relaxes to 0 at large times, confirming the stability of the localized solution found in Section V.2. The relaxation becomes very sluggish ($\nu_{\beta=0}\to 0^{-}$) when $\gamma$ or $q$ are small (i.e., $P_{0}^{(st)}$ small). In the curve $\gamma=0.4$ of Fig. 1b, one actually checks that the relaxation toward the steady state solution becomes faster as $q$ increases (see, e.g., the red circles).
@@
Therefore, when memory slowly decays with time, the relaxation toward the localized state is faster compared to the perfect memory case $\beta=0$.
$\bullet$ Case $\beta$ close to 11 : Let us set $\beta=1-\varepsilon_{1}$ and assume $\nu=-1+\varepsilon_{2}$, with $0<\varepsilon_{1}\ll 1$ and $0<\varepsilon_{2}\ll 1$. Using $\Gamma(1)=1$ and the expansion $\Gamma(z)\simeq 1/z$ at small $z$ in Eq. (53), one obtains $\varepsilon_{2}=\varepsilon_{1}/[(1-\gamma P_{0}^{(st)})^{-1}-1]$, or,
@@
Hence, when $\beta$ approaches 1 from below, the relaxation at large time can be much faster than for small values of $\beta$ and it follows a generic form approximately given by $h(t)\sim t^{-1}$, independently of $(q,\gamma)$. As shown below, this result suggests that the fastest relaxation toward the steady state, among the values of $\beta$ in the interval $[0,1]$, is achieved when $\beta$ is exactly 11.
$\bullet$ Boundary case $\beta=1$ : Eq. (53) no longer admits a solution and does not describe this case. After replacing the sum $\sum_{t^{\prime}=0}^{t}$ in Eq. (2) by $\int_{0}^{t}dt^{\prime}$, the memory kernel for $\beta=1$ reads $\pi_{t,t^{\prime}}\simeq[\ln(1+t)]^{-1}/(1+t-t^{\prime})$. Setting $\lambda=\lambda_{max}$ in the evolution equation (23) of $h(t)$ yields at large $t$,
@@
where once again h(t+1)h(t)h(t+1)-h(t) has been neglected compared to the two other terms. The power-law ansatz (50) of the relaxation at late times must be modified and we now assume
@@
The $1/t$ convergence toward the localized state thus has a logarithmic correction that depends on the parameters $(q,\gamma)$. If $\gamma P_{0}^{(st)}<1/2$, then $\alpha>0$ and the relaxation is a bit slower than for a stronger localization, i.e., $\gamma P_{0}^{(st)}>1/2$, where $\alpha<0$. When $\gamma P_{0}^{(st)}$ is close to 0, the exponent $\alpha$ becomes very large. Since $h(t)$ is a decreasing function of time, the asymptotic regime (58) with a large $\alpha$ is actually reached at extremely large times and very difficult to observe numerically.
@@
In Figure 6, we have represented the relaxation obtained by iterating Eq. (23) numerically for the leading mode $\lambda_{max}$, starting from the initial condition $h(t=0)=1$. As explained in the next subsection, we have displayed $h(t)/\ln t$ rather than $h(t)$ alone, as the former function includes the relaxation of the modes which are close to $\lambda_{max}$ instead of $\lambda_{max}$ alone. (Qualitatively similar results would be obtained by showing $h(t)$.) The figure also shows the quantity $P_{0}(t)-P_{0}^{(st)}$, where $P_{0}(t)$ is obtained from numerical simulations of a walker starting at the origin at $t=0$, while $P_{0}^{(st)}$ is given by Eq. (9). Although the agreement is not quantitative, the theory exhibits the correct shape for the relaxation. And, as predicted, the simulated relaxation becomes faster as $\beta$ increases from 0 and approaches 1 (yellow curve/dots). This effect is even stronger in the simulations than in theory.
@@ VI.2 Relaxation of the modes close to $k=0$
In the comparison with the numerical simulations of Fig. 6, we have also taken into account the eigenvalues that are close to $\lambda_{max}=1-q$ in the relaxation. As the eigenvalue spectrum $\lambda=(1-q)\cos k$ is continuous, the modes with $0<k\ll 1$ also contribute in a significant way and their overall effect turns out to be a logarithmic correction to the leading power-law decay (see [45] for a similar example). For any fixed $\beta\in[0,1)$, we find that the relaxation to the stationary state is given by
@@
where the dependence on $k$ has been made explicit in the functions $f_{n}$ and $h(t)$, while $a^{(i)}_{k}$ is to be determined from the initial condition. Applying the ansatz $h^{(k)}(t)=t^{\nu_{k}}$ to Eq. (23) with the eigenvalue $\lambda=(1-q)\cos k$ leads to a modified Eq. (53) for the exponent $\nu_{k}$ of that mode,
@@
with $c^{\prime}$ depending little on $k$ [recall that $h^{(k)}(t=0)=1$ by convention]. As clear from Eq. (64), at large $\ln t$ the integral in Eq. (61) is dominated by the small $k$ regime of a Gaussian and the integration domain can be extended to $k=\infty$. Fixing $n$ (e.g. $n=0$) one therefore needs to determine the behavior of $a_{k}^{(a,s)}$ and $f_{n}^{(a,s)(k)}$ at small $k$. Due to the cancellation of the double pole, $f_{n}^{(s)(k)}$ is dominated by the bound state (44) which depends little on $k$ close to $k=0$, i.e.,
@@
(up to prefactors) for the antisymmetric and symmetric solutions, respectively, with $\mu_{1}$ a function of $(q,\gamma)$. From Eq. (66), at fix $n$, one deduces that
@@
The coefficients $a_{k}^{(s)}$ follow from the initial perturbation $P_{n}(t=0)=P_{n}^{(st)}+g_{n}^{(0)}$ and from the orthogonality relation $\sum_{m=-\infty}^{\infty}\widetilde{f}_{m}^{(s)(k^{\prime})}f_{m}^{(s)(k)}=\delta(k-k^{\prime})$. Setting $h^{(k)}(t=0)=1$ in Eq. (61), multiplying by $\widetilde{f}_{n}^{(s)(k^{\prime})}$, summing over $n$ and integrating over $k^{\prime}$, one obtains
@@
In the last approximation of Eq. (68), the initial perturbation $g_{m}^{(0)}$ is assumed to be sufficiently localized near the origin. Gathering Eqs. (64), (65) and (68) into Eq. (61), one gets $P_{n}(t)-P_{n}^{(st)}\sim t^{-|\nu|}\int_{0}^{\infty}dk\,k\,e^{-ck^{2}\ln t}$, which yields Eq. (60) after integration.
@@ VII Dynamical scenarios of intermittent localization
From a dynamical system point of view, the stability analysis of Section V can be summarized by the phenomenological diagram of Fig. 7a, which sketches a high-dimensional space of distribution functions. Fixing the parameters $(q,\gamma)$ and assuming that memory is sufficiently long-range ($0\leq\beta\leq 1$), the purple fixed point represents the localized solution $P_{n}^{(st)}$ given by Eq. (10). All trajectories/initial conditions eventually converge to this point. We have singled out the one-dimensional manifold representing the eigenmode (44) with largest eigenvalue $\lambda_{max}$ or slowest decay.
@@
Based on the simulation results of Sections IV.2 and IV.3 for the exponential and power-law cases ($1<\beta\leq 2$), a possible qualitative picture for short-range memory systems clearly displaying intermittent localization is displayed in Fig. 7b. In this representation, we have assumed that the eigenstates are stable, except the one with the largest eigenvalue, which is unstable. This situation would arise, for instance, if the generating function has a single real pole with $|\epsilon_{1}|<1$ which is canceled by a sufficiently large $\lambda_{max}$ such that $\lambda_{max}>\lambda^{(I)}_{c}$. The stable manifold, on the other hand, must still be high-dimensional as it allows many trajectories to reach the vicinity of the localized state, although for a finite period of time due to the repulsion produced by the unstable manifold. The fixed point is essentially the same than in the stable case of Fig. 7a.
@@
(1) The system does not exhibit a transition between the cases of Fig. 7a and Fig. 7b for any parameter value $q$ or $\gamma$. In other words, it is $a)$ either always fully localized or $b)$ intermittently localized.
(2) The type of localization of a system can belong the cases of Fig. 7a or Fig. 7b, depending on $(q,\gamma)$. This means that there exists a critical value $q_{c}^{(I)}$ (which depends on $\gamma$ and the memory kernel parameters) such that the system is intermittently localized for $q<q_{c}^{(I)}$ and fully localized for $q>q_{c}^{(I)}$. This transition is produced by a change of stability of the most dangerous the eigenmode.
@@
We speculate that a necessary condition for intermittent localization is that the effective number of remembered visits $C(t)$ does not grow unbounded at late times but tends to a constant. This is the case for memory kernels that are exponential or algebraic with $\beta>1$. The numerical results clearly show that the mean localization time becomes extremely large as $q$ increases (Figs. 4a and 5c). However, at this point, we cannot determine precisely whether these examples belong to the case (1) or (2).

### v3

@@ V.2 Time-dependent part
The analysis of the spatial part has allowed us to find the largest eigenvalue (46) of the perturbation spectrum,
@@
In the temporal part, $\lambda$ only appears in the first term of the rhs of Eq. (26) for the evolution of $h(t)$, which involves the kernel $\pi_{t,t^{\prime}}$. We expect the amplitude to grow with time if $\lambda$ is large enough, say, above a threshold value denoted as $\lambda_{c}^{(I)}$, and to decay to 0 if $\lambda<\lambda_{c}^{(I)}$. The largest eigenvalue $\lambda_{max}$ corresponds to the most “dangerous” mode, which dominates the dynamics at late times by growing the fastest or decaying the slowest. Heuristically, the two types of asymptotic behaviors are separated by the marginal case at $\lambda=\lambda_{c}$ where $h(t)\to h^{*}$ at late times, with $h^{*}$ a constant. Substituting $h(t)$ by a constant in Eq. (26) and taking $t\to\infty$, one obtains
@@
For memory kernels of the form (2), where $F(\tau)$ is assumed to decay or to remain constant, the initial times in the interval $t^{\prime}\in[0,t_{0}]$ contribute very little at large $t$ to the distribution $\pi_{t,t^{\prime}}$. Hence, the sum in Eq. (49) is equal to unity due to normalization and the equation simplifies to
@@
These conclusions agree with some of our results. In Section IV.1, the fully localized states appeared to be stable attractors of the dynamics in the simulations, for sufficiently long-range memory kernels (see Fig. 1b). This is actually the regime where the de-correlation approximation used in the stability analysis is expected to be most accurate and valid.
Unfortunately, the same conclusion predicts that localization in short-range memory systems should be stable, too, which is inconsistent with the observation of intermittent localization. Therefore, the theory is unable to describe the intermittent localization phenomenon, for which one could have, for instance, at least one unstable mode or $\lambda_{max}>\lambda_{c}$. This suggests that during the evolution of localization toward diffusive states, temporal correlations start to play a crucial role and must be taken into account.
@@
Let us return to the case of the power-law memory kernel (18) with $0\leq\beta\leq 1$, a regime where our linear stability analysis holds in principle.
@@ VI.1 Relaxation of the slowest mode
The way the amplitude of a perturbation relaxes to zero does depend on the memory kernel. We can apply the results of the stability analysis to the study of the relaxation of the leading eigenmode with $\lambda=\lambda_{max}=1-q$, which corresponds to Eq. (43) with $k=0$. At large $t$, we approximate the sums by integrals and time is taken as continuous; in addition h(t+1)h(t)h(t+1)-h(t) is replaced by the time-derivative $\dot{h}(t)$. Equation (26) with $\lambda=1-q$ and $t_{0}=0$ becomes
@@
$h(t)=t^{\nu}\quad{\rm with}\quad\nu<0$ (53)
at late times, where the exponent $\nu$ is unknown. Following a similar method as in [34], we substitute $h(t^{\prime})$ by the form (53) over the whole time domain, choose a positive $\varepsilon\ll 1$ and a large $t$ such that $\varepsilon t\gg 1$. We then make the change $\tau=t-t^{\prime}$ and decompose,
@@
The first and second integrals in the rhs of Eq. (54) are ${\cal O}\left((\varepsilon t)^{1-\beta}\right)$ and ${\cal O}\left(\varepsilon(\varepsilon t)^{1-\beta}\right)$, respectively, and can be neglected compared to the third term, which is ${\cal O}\left(t^{1-\beta}\right)$. To see this, let us replace $(1+\tau)^{\beta}$ by $\tau^{\beta}$ in the denominator of the third term (since $\epsilon t\gg 1$),
@@
We analyze the different cases below.
$\bullet$ Case $\beta=0$ : When memory is perfect, the solution of Eq. (56) is simply $\nu_{\beta=0}=-\gamma P_{0}^{(st)}$, which belongs to the interval (1,0)(-1,0), as assumed. Hence,
@@
Hence, the amplitude of the most dangerous eigenmode relaxes to 0 at large times, confirming the stability of the localized solution found in Section V.2. The relaxation becomes very sluggish ($\nu_{\beta=0}\to 0^{-}$) when $\gamma$ or $q$ are small (i.e., $P_{0}^{(st)}$ small). In the curve $\gamma=0.4$ of Fig. 1b, one actually checks that the relaxation toward the steady state solution becomes slower as $q$ decreases (see, e.g., the red circles).
@@
Therefore, when memory slowly decays with time, the relaxation towards the localized state is faster compared to the perfect memory case $\beta=0$.
$\bullet$ Case $\beta$ close to $1$ : Let us set $\beta=1-\varepsilon_{1}$ and assume $\nu=-1+\varepsilon_{2}$, with $0<\varepsilon_{1}\ll 1$ and $0<\varepsilon_{2}\ll 1$. Using $\Gamma(1)=1$ and the expansion $\Gamma(z)\simeq 1/z$ at small $z$ in Eq. (56), one obtains $\varepsilon_{2}=\varepsilon_{1}/[(1-\gamma P_{0}^{(st)})^{-1}-1]$, or,
@@
Hence, when $\beta$ approaches 1 from below, the relaxation at large times is much faster than for small values of $\beta$ and it follows a generic form approximately given by $h(t)\sim t^{-1}$, independently of $(q,\gamma)$. As shown below, this result suggests that the fastest relaxation toward the steady state, among the values of $\beta$ in the interval $[0,1]$, is achieved when $\beta$ is exactly $1$.
$\bullet$ Boundary case $\beta=1$ : Eq. (56) no longer admits a solution and does not describe this case. After replacing the sum $\sum_{t^{\prime}=0}^{t}$ in Eq. (2) by $\int_{0}^{t}dt^{\prime}$, the memory kernel for $\beta=1$ can be approximated by $\pi_{t,t^{\prime}}\simeq[\ln(1+t)]^{-1}/(1+t-t^{\prime})$. Setting $\lambda=\lambda_{max}$ in the evolution equation (26) of $h(t)$ yields at large $t$,
@@
where once again h(t+1)h(t)h(t+1)-h(t) has been neglected compared to the other two terms. The power-law ansatz (53) of the relaxation at late times must be modified and we now assume
@@
The $1/t$ convergence toward the localized state exhibits a logarithmic correction that depends on the parameters $(q,\gamma)$. If $\gamma P_{0}^{(st)}<1/2$, then $\alpha>0$ and the relaxation is a bit slower than for a stronger localization, i.e., $\gamma P_{0}^{(st)}>1/2$, where $\alpha<0$. When $\gamma P_{0}^{(st)}$ is close to 0, the exponent $\alpha$ becomes very large. Since $h(t)$ is a decreasing function of time, the asymptotic regime (61) with a large $\alpha$ is actually reached at extremely large times and is very difficult to observe numerically.
@@
In Figure 6, we have represented the relaxation obtained by iterating Eq. (26) numerically for the leading mode $\lambda_{max}$, starting from the initial condition $h(t=0)=1$ (solid lines). As explained in the next subsection, we have displayed $h(t)/\ln t$ rather than $h(t)$ alone, as the former function includes the relaxation of the modes which are close to $\lambda_{max}$ instead of $\lambda_{max}$ alone. [Qualitatively similar results are obtained by representing $h(t)$.] The figure also shows the quantity $P_{0}(t)-P_{0}^{(st)}$, where $P_{0}(t)$ is obtained from numerical simulations of a walker starting at the origin at $t=0$, while $P_{0}^{(st)}$ is given by Eq. (12). Although the agreement is not quantitative, the theory exhibits the correct relaxation shape. And, as predicted, the simulated relaxation becomes faster as $\beta$ increases from 0 and approaches 1 (yellow curve/dots). This effect is even stronger in the simulations than in theory.
@@ VI.2 Relaxation of the modes close to $k=0$
In the comparison with the numerical simulations of Fig. 6, we have also taken into account the eigenvalues that are close to $\lambda_{max}=1-q$ in the relaxation. As the eigenvalue spectrum $\lambda=(1-q)\cos k$ is continuous, the modes with $0<k\ll 1$ significantly contribute, too, and their overall effect turns out to be a logarithmic correction to the leading power-law decay (see [49] for a similar example). For any fixed $\beta\in[0,1)$, we find that the relaxation to the stationary state is given by
@@
where the dependence on $k$ has been made explicit in the functions $f_{n}$ and $h(t)$, while the coefficients $a^{(i)}_{k}$ is to be determined from the initial condition. Applying the ansatz $h^{(k)}(t)=t^{\nu_{k}}$ to Eq. (26) with the eigenvalue $\lambda=(1-q)\cos k$ leads to a modified Eq. (56) for the exponent $\nu_{k}$ of that mode,
@@
with $c^{\prime}$ depending little on $k$ [recall that $h^{(k)}(t=0)=1$ by convention]. As clear from Eq. (67), at large $\ln t$ the integral in Eq. (64) is dominated by the small $k$ regime of a Gaussian and the integration domain can be extended to $k=\infty$. Fixing $n$ (e.g. $n=0$) we therefore need to determine the behavior of $a_{k}^{(a,s)}$ and $f_{n}^{(a,s)(k)}$ at small $k$. Due to the cancellation of the double pole, $f_{n}^{(s)(k)}$ is dominated by the bound state (47) which depends little on $k$ close to $k=0$, i.e.,
@@
(up to prefactors) for the antisymmetric and symmetric solutions, respectively, with $\mu_{1}$ a function of $(q,\gamma)$. From Eq. (69), at fixed $n$, one deduces the small $k$ behavior
@@
The coefficients $a_{k}^{(s)}$ follow from the initial perturbation $P_{n}(t=0)=P_{n}^{(st)}+g_{n}^{(0)}$ and from the orthogonality relation $\sum_{m=-\infty}^{\infty}\widetilde{f}_{m}^{(s)(k^{\prime})}f_{m}^{(s)(k)}=\delta(k-k^{\prime})$. Setting $h^{(k)}(t=0)=1$ in Eq. (64), multiplying this equation by $\widetilde{f}_{n}^{(s)(k^{\prime})}$, summing over $n$ and integrating over $k^{\prime}$, one obtains
@@
In the last approximation of Eq. (71), the initial perturbation $g_{m}^{(0)}$ is assumed to be sufficiently localized near the origin (i.e., vanishing for $|m|\gg 1$). Gathering Eqs. (67), (68) and (71) into Eq. (64), one gets $P_{n}(t)-P_{n}^{(st)}\sim t^{-|\nu|}\int_{0}^{\infty}dk\,k\,e^{-ck^{2}\ln t}$, which yields Eq. (63) after integration.
@@ VII Dynamical scenarios of intermittent localization
From a dynamical system point of view, the stability analysis of Section V can be summarized by the phenomenological diagram of Fig. 7a, which sketches a high-dimensional space of distribution functions. Fixing the parameters $(q,\gamma)$ and assuming that memory is sufficiently long-range ($0\leq\beta\leq 1$), the purple fixed point represents the localized solution $P_{n}^{(st)}$ given by Eq. (13). All trajectories/initial conditions eventually converge to this point. We have singled out the one-dimensional manifold representing the eigenmode (47) with the largest eigenvalue $\lambda_{max}$ or slowest decay.
@@
Based on the simulation results of Sections IV.2 and IV.3 for the exponential and power-law cases ($1<\beta\leq 2$), a possible qualitative picture for short-range memory systems is displayed in Fig. 7b. In this representation, we have assumed that the eigenstates are stable, except the one with the largest eigenvalue, which is unstable. This situation would arise, for instance, if the generating function had a single real pole with $|\epsilon_{1}|<1$ which is canceled by a sufficiently large $\lambda_{max}$, such that $\lambda_{max}>\lambda^{(I)}_{c}$. The stable manifold, on the other hand, must still be high-dimensional as it allows many trajectories to reach the vicinity of the localized state, although for a finite period of time due to the repulsion produced by the unstable manifold. The fixed point is essentially the same as in the stable case of Fig. 7a.
@@
(1) The system does not exhibit a transition between the cases of Fig. 7a and Fig. 7b for any parameter value $q$ or $\gamma$. In other words, it is either $a)$ always fully localized, or $b)$ always intermittently localized.
(2) The type of localization of a system can belong to the cases of Fig. 7a or Fig. 7b, depending on $(q,\gamma)$. This means that there exists a critical value $q_{c}^{(I)}$ (which depends on $\gamma$ and the memory kernel parameters) such that the system is intermittently localized if $q<q_{c}^{(I)}$ and fully localized if $q>q_{c}^{(I)}$. This transition is produced by a change in stability of the most dangerous eigenmode.
@@
We speculate that a necessary condition for intermittent localization is that the effective number of remembered visits $C(t)$ does not grow unbounded at late times but tends to a constant. This is the case for memory kernels that are exponential or algebraic with $\beta>1$. The numerical results clearly show that the mean localization time becomes extremely large as $q$ increases (Figs. 4a and 5c). However, at this point, we cannot determine precisely whether these examples belong to the cases (1) or (2).